OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 + (sin Θ + cos Θ)2

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@mtbender74

OpenStudy (kainui):

Well what have you tried? Looks like you should probably square these and see if anything simplifies. Also don't forget the pythagorean identity, \[\sin^2 \theta + \cos^2 \theta = 1\]

OpenStudy (hba):

As @Kainui said you should expand it and use this formula.

OpenStudy (anonymous):

ok :)